4. A decision at the margin Hilary is a hard-working college senior. One Thursday, she decides to work nonstop until she has answered 100 practice problems for her chemistry course. She starts work at 8:00 AM and uses a table to keep track of her progress throughout the day. She notices that as she gets tired, it takes her longer to solve each problem. Time 8:00 AM 9:00 AM 10:00 AM 11:00 AM Total Problems Answered 40 70 90 100 Noon Use the table to answer the following questions The marginal, or additional, gain from Hilarys first hour of work, from 8:00 AM to 9:00 AM, 40 problems. The marginal gain from Hilarys third hour of work, from 10:00 AM to 11:00 AM, Is 20 problems Later, the teaching assistant in Hilarys chemistry course gives her some advice. Based on past experience, the teaching assistant says, working on 35 problems raises a students exam score by about the same amount as reading the textbook for 1 hour. For simplicity, assume students always cover the same number of pages during each hour they spend reading. Given this information, in order to use her 4 hours of study time to get the best exam score possible, how many hours should she have spent working on problems, and how many should she have spent reading?
mework (Ch 01) 10:00 AM 11:00 AM Noon 70 90 100 Use the table to answer the following questions The marginal, or additional, gain from Hilarys first hour of work, from 8:00 AM to 9:00 AM, is 40 problems The marginal gain from Hilarys third hour of work, from 10:00 AM to 11:00 AM, is 20 problems Later, the teaching assistant in Hilarys chemistry course gives her some advice. Based on past experience, the teaching assistant says, working on 35 problems raises a students exam score by about the same amount as reading the textbook for 1 hour.For simplicity, assume students always cover the same number of pages during each hour they spend reading. Given this information, in order to use her 4 hours of study time to get the best exam score possible, how many hours should she have spent working on problems, and how many should she have spent reading? 0 hours working on problems, 4 hours reading O 1 hour working on problems, 3 hours reading 3 hours working on problems, 1 hour reading 4 hours working on problems, O hours reading Grade it Now Save & Continue Continue without saving
Answer #1

Marginal gain is the additional gain by solving problems for one more hour.

1) Marginal gain from first hour = Problems solved from 9 to 10 - Problems solved from 8 to 9 = 40 - 0 = 40 problems

2) Marginal gain from third hour = 90 - 70 = 20 problems

3) 1 hour working on problems, 3 hours reading

Working on problems for 1 hours gives student score of 40 while another hour on problems does not increase student score by more than 30 so student should prefer reading which increases score by 35.